Golem:

Does anyone know if an exponent of -1 to any number just changes the number into a negative or something else

THEKINGOFTHEMATRIX:

Probably a positive I don't exactly know for sure bro ¯\_(ツ)_/¯

carmelle:

The number actually becomes a fraction/decimal... Example: \[10^{-1} \rightarrow \frac{ 1 }{ 10 }\] \[30^{-1} \rightarrow \frac{ 1 }{ 30 }\] However, if the number is negative, then the fraction will be negative... Example: \[-30^{-1} \rightarrow -\frac{ 1 }{ 30 }\] \[-40^{-1} \rightarrow -\frac{ 1 }{ 40 }\]

Joe348:

but if a number was \[-1^2\] your answer would be i